Find the static thrust of a turbojet engine on a test stand at sea level standard day conditions in Newtons and then convert to lbf. Assume mass flow rate is constant.

Inlet and Exit areas are both equal to 0.75 m2

Exit velocity = 600 m/s

Exit pressure = 1.0 atm

Exit temperature = 725 K
Answer #1

First of all, we will find the mass flow,

\rhoe = Pe / RTe

\rhoe = 1.01e5 / 287 * 725

\rhoe = 0.4854 kg / m3

Therefore, mass flow rate is

m = \rho e Aeve

m =  0.4854 *0.75 * 600

m = 218.43 kg/sec

Now, velocity at the inlet

vi = 218.43 / 1.225 * 0.75

where 1.225 is the density at sea level

vi = 237.74 m/s

Now, I will use the thrust equation

T = m (ve - vi) + (pe - p)Ae

T = m (ve - vi)

T = 218.43 (600 - 237.74)

T = 79126.94 N

Now, we know

1 lb = 4.448 N

therefore,

T = 17789.7 lb